OpenStudy (anonymous):

Please help :) Solve the equation: 87+2x/216+3x

OpenStudy (anonymous):

do you mean = [87+(2x /216)+3x\]

OpenStudy (anonymous):

=0?

OpenStudy (anonymous):

To be an equation, you need an = sign

OpenStudy (anonymous):

=.5

OpenStudy (anonymous):

Go on, noru...do it.

OpenStudy (anonymous):

if 87 + (2x/216) + 3x = 0.5 x ((2/16)+3)= -86.5 x= - 86.5 * 3.00925 x= - 260.3

OpenStudy (anonymous):

2/16?

OpenStudy (anonymous):

Lemme check...

OpenStudy (anonymous):

sorry 2/216